OpenStudy (anonymous):

Sorry, wrong question! the integral of (x(ln(1+x))dx

OpenStudy (zarkon):

integration by parts

OpenStudy (anonymous):

xdx=(x^2)/2+c

OpenStudy (zarkon):

?

OpenStudy (zarkon):

have you done integration by parts before?

OpenStudy (anonymous):

Yes, but I cannot figure this one out. What shoul u be? ln(1+x)?

OpenStudy (zarkon):

\[u=\ln(1+x)\] \[dv=xdx\]

OpenStudy (anonymous):

Used that, but just keeps growing.

OpenStudy (zarkon):

how ...

OpenStudy (zarkon):

\[du=\frac{1}{1+x}dx\] \[v=\frac{x^2}{2}\] \[\int x\ln(1+x)dx=\frac{x^2}{2}\ln(1+x)-\int\frac{x^2}{2(1+x)}dx\]

OpenStudy (zarkon):

right?

OpenStudy (zarkon):

a little algebra gives us \[\frac{x^2}{2(1+x)}=\frac{1}{2(1+x)}+\frac{x}{2}-\frac{1}{2}\]

OpenStudy (anonymous):

partial fractions?

OpenStudy (zarkon):

long division

OpenStudy (zarkon):

you don't use partial fractions when the power if the numerator is >= power of the denominator

OpenStudy (anonymous):

thanks got it now

OpenStudy (zarkon):

good